# JWST provides a new view of cosmic dawn: latest developments in studies of early galaxies

### v1

@@ 4.1.3 Rapid formation and quenching
Of particular interest are galaxies with Balmer breaks in their spectra, as those indicate that relatively old stellar populations already exist (Figure 3). Since their star formation has quenched for a while, these galaxies are also used to study the causes of quenching. While passive galaxies had already been spectroscopically confirmed at before JWST [e.g. Schreiber18], JWST spectroscopy allows the measurement of faint continuum and absorption line features with far higher precision and out to higher redshifts [e.g. Carnall23]. The sensitive photometric band coverage by NIRCam is improving our ability to perform statistical analyses and obtain robust estimates of their number densities [Valentino23], which are needed to compare to models of structure formation.
@@ 4.2.1 The reionization timeline
The progress of reionization can also be mapped using observed statistics of the emerging Ly emission-line from galaxies, such as their escape fraction and line-profile [e.g. Hayes23]. The observed Ly line profiles are similarly impacted by the Ly damping wing and are more easily observed than faint continuum features. Early estimates of the evolution of the neutral fraction have been made using Ly demographics of distant galaxies [Tang24, Kageura25]. Various studies report a connection between large-scale over-densities and the observability of Ly emission, effectively extending the connection between galaxy over-densities and ionized regions out to [e.g. Saxena23, Whitler24, Witstok24]. Current efforts are pushing these measurements to the highest redshifts [Witstok25] and aiming at separating the relative impacts of bright and faint galaxies in their connection to ionised bubbles [e.g. Torralba24]. Moreover, as the Ly escape fraction and line-profile are also strongly sensitive to the gas conditions of galaxies, a detailed understanding of the correlations among those properties is required to make Ly a reliable tracer of the progress of reionization [e.g. Lin24].
@@ 4.3.1 Metallicity and abundance measurements
The study of gas-phase metal abundances in interstellar gas has a long historical record [Tinsley79, Tremonti04, Andrews13] and was reviewed recently by Curti [Curti25]. The variation of metal abundances within and among galaxies depends on the star formation history and the nucleosynthetic origins of the elements, but also stellar winds and the accretion of newly accreted material [e.g. Kobayashi25]. The emission-lines that have traditionally been used to measure the gas-phase metallicity of galaxies are strong lines in the rest-frame optical. These strong lines trace common metals in the Universe (oxygen, neon, nitrogen) and are usually observed together with hydrogen emission lines from the Balmer series. Since oxygen is the most abundant metal and usually the strongest metal emission-line, the majority of gas-phase abundances are oxygen abundance measurements, especially at high-redshift. This is an important distinction from measurements of metallicities in stars, which usually are iron abundances [Steidel16, Strom22]. Significant challenges exist in the measurement of gas-phase metallicities. Emission-line ratios are also sensitive to other properties of the gas, such as the ionization parameter, electron temperature, and electron density.
@@ 4.3.2 The first stars and galaxies: expectations and early insights
How would one identify a PopIII galaxy (or a clump of PopIII stars within a galaxy)? The cleanest detection would be the total absence of metals in the spectrum of a star/stellar population 999In the Milky Way, stars with metallicities as low as [Z/H] have been reported [Caffau11]., however, this would require extremely sensitive spectroscopy beyond our current capabilities (see [Schauer20] for a discussion). More feasible strategies have either focused on identifying a very low gas-phase metallicity (mainly constrained through the oxygen abundance; e.g. [Trussler23, Fujimoto25]) or on identifying expected signatures of PopIII stars such as very blue UV continua, unusual continuum shapes and strong HeII line emission. The latter properties are in some way sensitive to the predicted hot temperatures of PopIII stars and its impact on nebular emission [Zackrisson24]. The challenges associated with these techniques are that PopIII stars could be missed if the surrounding gas has been enriched from external sources, metal line emission could be quenched due to high densities (e.g. in case of the [Oiii]5008 line) and that other ionising sources may mimic PopIII spectral signatures (i.e. HeII can be powered by PopII stars stripped in binary interactions, or by AGN, shocks [Lecroq24]).
@@ 4.4.1 Quasars: properties, environments and hosts
JWST observations of quasars have focused on probing the stellar masses of the systems hosting quasars. Despite the high resolution imaging data that JWST can provide, detecting host galaxies is extremely challenging due to outshining by quasars. For example, a typical quasar at has an AB magnitude of at 2 micron, whereas the host galaxy would have a magnitude of for a stellar mass of M⊙. Careful subtraction of the quasar emission has revealed various indications of extended massive host galaxy detections (with masses M⊙; [e.g. Yue24]), but in some cases showing complex morphologies that indicate merger activity [e.g. Marshall24]. Spectroscopic studies of somewhat fainter quasars have directly identified stellar emission from the host galaxy in the rest-frame optical [Ding23, Onoue24], yielding more precise estimates of their stellar masses. Generally, despite the high inferred stellar masses, the BH to stellar mass ratios are on the order %, demonstrating the efficient formation of SMBHs in the early Universe compared to the assembly of stellar mass.
Independent insights into the systems that host quasars can be obtained from their large-scale clustering. Given their extremely low number densities ( cMpc-3; [Schindler23]), one would naively expect that they are hosted by rare massive dark matter halos with masses M⊙. As the number densities of quasars are too low to perform quasar-quasar clustering measurements, the halo masses of quasars are inferred through their correlation with galaxies in the environment. JWSTs NIRcam wide field slitless spectroscopic mode is highly efficient at mapping the large-scale environments on cMpc scales in the plane of the sky [e.g. Kashino23], see Section 3.3. The average quasar resides in relatively large over-densities of galaxies, however, there appears to be a large variation in the galaxy over-densities around high-redshift quasars [FWang23, Eilers24]. The average dark matter halo mass implied from the galaxy over-densities is M⊙, which is similar to quasars seen at later epochs, i.e. , [Laurent17]. This is well below the halo mass that would correspond to the observed number densities of quasars. This implies that luminous quasars at do not reside in the most massive halos that exist at that cosmic time, but also that the duty cycle, which is the fraction of such halos that are detectable as UV-luminous quasars at a given epoch, is low.
@@ 4.4.2 Faint AGN & little red dots
Thus, it seems that the LRDs are revealing a previously unknown phase of gas-enshrouded AGN formation that is most prevalent in the early Universe [Inayoshi25, Pacucci25]. However, many things remain unclear. What is the origin of the dense gas? Is it an envelope being expelled from an early ‘quasi-star’ phase that nurtures the rapid formation of supermassive black holes [Begelman25] or is the high covering fraction an indicator of near-spherical accretion flows theoretically expected for the most rapid accretion regimes [Liu25]? What are the masses of the black holes powering LRD emission? Are there nearby star clusters associated to these black holes? Which mechanisms facilitate the rapid formation of massive black holes and how does the growth of these black holes impact their host galaxy? With the rapid advances being made in this field, one thing that is clear is that we are about to learn something new.
@@ 5 Outlook
The JWST has been operational for only three years, but it has revolutionized our understanding of the early Universe. Its performance has been exceeding expectations [Rigby23], but the Universe itself has also kept surprising astronomers. JWST data have highlighted new discovery spaces that can be applied to other data-sets. For example, the prominent nitrogen lines in the rest-frame UV spectrum of GN-z11 at have motivated investigations into the abundances and gas densities in galaxies and globular clusters in the local Universe. JWST has unveiled a previously unknown population of faint AGN with highly unusual spectra and physical conditions. In retrospect, such AGNs appeared already present at in the SDSS and were studied in 2008 [Izotov08], but were not given that much attention because they are very rare in the local Universe and earlier spectra did not identify unusual features as the Balmer line absorption. It may also well be that various seemingly separate developments may turn out to be quite related. For example, nitrogen-enriched gas has been associated to (nuclear) star cluster formation that may be the sites in which early SMBH growth happens [e.g. Isobe25], whereas the mechanisms that cause bursty star formation in early galaxies could also disturb equilibrium scaling relations in variations among metallicity and star formation rates [e.g. McClymont25NO].
In the next few years, detailed JWST follow-up studies will target the newly identified populations of galaxies, while prioritising wavelength ranges of special interest with very deep spectroscopy. Simultaneously, wide-area surveys that are starting or about to start, such as the optical Legacy Survey of Space and Time by the Vera C. Rubin Observatory and the sensitive wide-area survey in the near-infrared by the Euclid and Roman telescopes and their spectroscopic follow-up campaigns, are poised to identify new exciting targets such as low-redshift LRDs or very high-redshift quasars. The Extremely Large Telescope (ELT) is also imminent, with first light currently planned by the end of this decade. Due to its large light gathering power and the possibility to design more complicated instruments, we can primarily expect significantly improved (rest-frame UV) spectra of galaxies in the distant Universe. The Square Kilometre Array may detect the 21-cm signal from the dark ages, which will enable new cosmological and astrophysical constraints, particularly on the sources and timeline of reionization. Finally, the lessons learned from these observational facilities are being incorporated in new, more complete models of galaxy formation.
@@ Acknowledgements
I thank Claudia Di Cesare, Edoardo Iani, Gauri Kotiwale and Wendy Sun for proofreading, Daichi Kashino, Gauri Kotiwale, Sara Mascia, Benjamín Navarette and Joris Witstok for their assistance in preparing some of the Figures, and Richard Ellis and Stephen Blundell for constructive comments.

### v2

@@ 4.1.3 Rapid formation and quenching
Of particular interest are galaxies with Balmer breaks in their spectra, as those indicate that relatively old stellar populations already existed (Figure 3). Since their star formation has quenched for a while, these galaxies are also used to study the causes of quenching. While passive galaxies had already been spectroscopically confirmed at before JWST [e.g. Schreiber18], JWST spectroscopy allows the measurement of faint continuum and absorption line features with far higher precision and out to higher redshifts [e.g. Carnall23]. The sensitive photometric band coverage by NIRCam is improving our ability to perform statistical analyses and obtain robust estimates of their number densities [Valentino23], which are needed to compare to models of structure formation.
@@ 4.2.1 The reionization timeline
The progress of reionization can also be mapped using observed statistics of the emerging Ly emission-line from galaxies, such as their escape fraction and line-profile [e.g. Hayes23]. The observed Ly line profiles are similarly impacted by the Ly damping wing and are more easily observed than faint continuum features. Early estimates of the evolution of the neutral fraction have been made using Ly demographics of distant galaxies [Tang24, Kageura25]. Various studies report a connection between large-scale over-densities and the observability of Ly emission, effectively extending the connection between galaxy over-densities and ionized regions out to [e.g. Saxena23, Whitler24, Witstok24]. Current efforts are pushing these measurements to the highest redshifts [Witstok25] and aiming at separating the relative impacts of bright and faint galaxies in their connection to ionised bubbles [e.g. Torralba24]. Moreover, as the Ly escape fraction and line-profile are also strongly sensitive to the interstellar gas conditions in galaxies, a detailed understanding of the correlations among those properties is required to make Ly a reliable tracer of the progress of reionization [e.g. Lin24].
@@ 4.3.1 Metallicity and abundance measurements
The study of gas-phase metal abundances in interstellar gas has a long historical record [Tinsley79, Tremonti04, Andrews13] and was reviewed recently by Curti [Curti25]. The variation of metal abundances within and among galaxies depends on the star formation history and the nucleosynthetic origins of the elements, but also on stellar winds and the accretion of newly accreted material [e.g. Kobayashi25]. The emission-lines that have traditionally been used to measure the gas-phase metallicity of galaxies are strong lines in the rest-frame optical. These strong lines trace common metals in the Universe (oxygen, neon, nitrogen) and are usually observed together with hydrogen emission lines from the Balmer series. Since oxygen is the most abundant metal and usually the strongest metal emission-line, the majority of gas-phase abundances are oxygen abundance measurements, especially at high redshift. This is an important distinction from measurements of metallicities in stars, which usually are iron abundances [Steidel16, Strom22]. Significant challenges exist in the measurement of gas-phase metallicities because emission-line ratios are also sensitive to other properties of the gas, such as the ionization parameter, electron temperature, and electron density.
@@ 4.3.2 The first stars and galaxies: expectations and early insights
How would one identify a Pop III galaxy (or a clump of Pop III stars within a galaxy)? The cleanest detection would be the total absence of metals in the spectrum of a star/stellar population 999In the Milky Way, stars with metallicities as low as [Z/H] have been reported [Caffau11]., however, this would require extremely sensitive spectroscopy beyond our current capabilities (see [Schauer20] for a discussion). More feasible strategies have either focused on identifying a very low gas-phase metallicity (mainly constrained through the oxygen abundance; e.g. [Trussler23, Fujimoto25]) or on identifying expected signatures of Pop III stars such as very blue UV continua, unusual continuum shapes and strong Heii line emission. The latter properties are in some way sensitive to the predicted hot temperatures of Pop III stars and its impact on nebular emission [Zackrisson24]. The challenges associated with these techniques are that Pop III stars could be missed if the surrounding gas has been enriched from external sources, metal line emission could be quenched due to high densities (e.g. in case of the [Oiii]5008 line) and that other ionising sources may mimic PopIII spectral signatures (i.e. Heii can be powered by Pop II stars stripped in binary interactions [Gotberg19], or by AGN, shocks [Lecroq24]).
@@ 4.4.1 Quasars: properties, environments and hosts
JWST observations of quasars have focused on probing the stellar masses of the systems hosting quasars. Despite the high resolution imaging data that JWST can provide, detecting host galaxies is extremely challenging due to outshining by quasars. For example, a typical quasar at has an AB magnitude of at 2 micron, whereas the host galaxy would have a magnitude of for a stellar mass of M⊙. Careful subtraction of the quasar emission has revealed various indications of extended massive host galaxy detections (with masses M⊙; [e.g. Yue24]), but in some cases showing complex morphologies that indicate merger activity [e.g. Marshall24]. Spectroscopic studies of somewhat fainter quasars have directly identified stellar emission from the host galaxy in the rest-frame optical [Ding23, Onoue24], yielding more precise estimates of their stellar masses. Generally, despite the high inferred stellar masses, the BH to stellar mass ratios are on the order of %, demonstrating the efficient formation of SMBHs in the early Universe compared to the assembly of stellar mass.
Independent insights into the systems that host quasars can be obtained from their large-scale clustering. Given their extremely low number densities ( cMpc-3; [Schindler23]), one would naively expect that they are hosted by rare massive dark matter halos with masses M⊙. As the number densities of quasars are too low to perform quasar-quasar clustering measurements, the halo masses of quasars are inferred through their correlation with galaxies in the environment. JWST’s NIRcam wide field slitless spectroscopic mode is highly efficient at mapping the large-scale environments on cMpc scales in the plane of the sky [e.g. Kashino23], see Section 3.3. The average quasar resides in relatively large over-densities of galaxies, however, there appears to be a large variation in the galaxy over-densities around high-redshift quasars [FWang23, Eilers24]. The average dark matter halo mass implied from the galaxy over-densities is M⊙, which is similar to quasars seen at later epochs, i.e. , [Laurent17]. This is well below the halo mass that would correspond to the observed number densities of quasars. This implies that luminous quasars at do not reside in the most massive halos that exist at that cosmic time, but also that the duty cycle, which is the fraction of such halos that are detectable as UV-luminous quasars at a given epoch, is low.
@@ 4.4.2 Faint AGN & little red dots
Thus, it seems that the LRDs are revealing a previously unknown phase of gas-enshrouded AGN formation that is most prevalent in the early Universe [Inayoshi25, Pacucci25]. However, many things remain unclear. What is the origin of the dense gas? Is it an envelope being expelled from an early ‘quasi-star’ phase that nurtures the rapid formation of supermassive black holes [Begelman25] or is the high covering fraction an indicator of near-spherical accretion flows theoretically expected for the most rapid accretion regimes [Liu25]? What are the masses of the black holes powering LRD emission? Are there nearby star clusters associated to these black holes? Which mechanisms facilitate the rapid formation of massive black holes and how does the growth of these black holes impact their host galaxy? With the rapid advances being made in this field, one thing is clear: we are about to learn something new.
@@ 5 Outlook
The JWST has been operational for only three years, but it has revolutionized our understanding of the early Universe. Its performance has been exceeding expectations [Rigby23], but the Universe itself has also kept surprising astronomers. JWST data have highlighted new discovery spaces that can be applied to other data-sets. For example, the prominent nitrogen lines in the rest-frame UV spectrum of GN-z11 at have motivated investigations into the abundances and gas densities in galaxies and globular clusters in the local Universe. JWST has unveiled a previously unknown population of faint AGN with highly unusual spectra and physical conditions. In retrospect, such AGNs appeared already present at in the SDSS and were studied in 2008 [Izotov08], but were not given that much attention because they are very rare in the local Universe and earlier spectra did not identify unusual features as the Balmer line absorption. It may also well be that various seemingly separate developments may turn out to be quite related. For example, nitrogen-enriched gas has been associated with (nuclear) star cluster formation that may be the sites in which early SMBH growth happens [e.g. Isobe25], whereas the mechanisms that cause bursty star formation in early galaxies could also disturb equilibrium scaling relations in variations among metallicity and star formation rates [e.g. McClymont25NO].
In the next few years, detailed JWST follow-up studies will target the newly identified populations of galaxies, while prioritising wavelength ranges of special interest with very deep spectroscopy. Simultaneously, wide-area surveys that are starting or about to start, such as the optical Legacy Survey of Space and Time by the Vera C. Rubin Observatory and the sensitive wide-area survey in the near-infrared by the Euclid and Roman telescopes and their spectroscopic follow-up campaigns, are poised to identify new exciting targets such as low-redshift LRDs or very high-redshift quasars. The Extremely Large Telescope is also imminent, with first light currently planned by the end of this decade. Due to its large light gathering power and the possibility to design more complicated instruments, we can primarily expect significantly improved (rest-frame UV) spectra of galaxies in the distant Universe and images at higher spectral and spatial resolution. The Square Kilometre Array may detect the 21-cm signal from the dark ages, which will enable new cosmological and astrophysical constraints, particularly on the sources and timeline of reionization. Finally, the lessons learned from these observational facilities are being incorporated in new, more complete models of galaxy formation.
@@ Acknowledgements
I thank Claudia Di Cesare, Edoardo Iani, Gauri Kotiwale and Wendy Sun for proofreading, Daichi Kashino, Gauri Kotiwale, Sara Mascia, Benjamín Navarrete and Joris Witstok for their assistance in preparing some of the Figures, and Richard Ellis and Stephen Blundell for constructive comments.